A particle travels at 1.90×108 m/s and lives 2.10×10−8 s when at rest relative to an observer. How long does the particle live as viewed in the laboratory?

c = 3.0×108

1
Expert's answer
2022-04-24T17:19:46-0400

"t=\frac{t_0}{\sqrt{1-\frac{v^2}{c^2}}}=2.7\cdot 10^{-8}~s."